OpenStudy (anonymous):

prove that the directrix of an ellipse is \[x=\frac{a^2}{c}\]

OpenStudy (anonymous):

this ellipse is the general ellipse with \[\frac{ x^2 }{ a^2}+\frac{ y^2 }{ b^2 }=1,a>b\]

OpenStudy (anonymous):

i have the proof but i dont understand the ratio set up here can i post the link

OpenStudy (anonymous):

http://mathworld.wolfram.com/Ellipse.html

OpenStudy (anonymous):

\[r=\frac{ a-c }{ d-a }=\frac{ a }{ d }\]

OpenStudy (anonymous):

\[a^2=b^2+c^2 i dont get this ratio .why it holds

OpenStudy (experimentx):

??